OpenStudy (anonymous):

Help will Medal In a race, a bicyclist travels 2/3 of a mile in 1/3 hour. Determine the unit rate for the speed in miles per hour. 1/2 mile per hour 1 mile per hour 4/3 miles per hour 2 miles per hour

OpenStudy (anonymous):

speed = distance / time speed = (2/3)/ (1/3) speed = 2*3/3 speed = 2 miles per hour

OpenStudy (anonymous):

thnx